# Tournament design: A review from an operational research perspective

### v1

@@ Abstract
Every sport needs rules. Tournament design refers to the rules that determine how a tournament, a series of games between a number of competitors, is organized. This study aims to provide an overview of the tournament design literature from the perspective of operational research. Three important design criteria are discussed: efficacy and effectivity, fairness, and attractiveness. Our survey classifies the papers discussing these properties according to the main components of tournament design: format, seeding, draw, scheduling, and ranking. We also outline several open questions and promising directions for future research.
@@ 1 Introduction
Practicing sports is an important part of the life of many people all over the world as many health, psychological, and social benefits are associated with physical activity. Watching and following sports is even more popular than playing it. About 3 billion people watched the 2020 Summer Olympic Games (Statista,, 2023), while no less than 1.5 billion people tuned in for the 2022 FIFA World Cup final (FIFA, 2023b, ). Even tournament design aspects can be popular: the draw for the 2020 UEFA Champions League Round of 16 has attracted more than 1.3 million views (Boczoń and Wilson,, 2023). Hence, “designing an optimal contest is both a matter of significant financial concern for the organizers, participating individuals, and teams, and a matter of consuming personal interest for millions of fans(Szymanski,, 2003, p. 1137).
@@
Every tournament is governed by a set of rules that determine the set of players, the format, the schedule, the ranking, and the prize allocation. Deciding on these rules is what we call tournament design. The set of matches to be played is given by the tournament format. Seeding is typically used to guarantee that a strong contestant is not eliminated at the early stages of the tournament merely because it plays against an even stronger opponent. Subject to these and other constraints required by the organizer, the contestants can be allocated randomly, through a so-called draw. At first sight, the schedule, the order and timing of the matches, seems a purely operational issue (involving travel distance, referee and venue availability, etc.). However, it may have powerful effects on the fairness and the outcome of the tournament. Ranking aims to order the teams according to their performance in the tournament. Although performance can be directly measured in some sports (e.g. by finishing times), most sports – especially those contested by teams – allow only to rank the contestants by paired comparisons, games or matches. Prize allocation determines how prizes (e.g. promotion, relegation, qualification, broadcasting revenue) will be awarded to the contestants, partially depending on the ranking.
@@
Given the variety of tournament designs, the question arises on what basis they can be compared. We identify four different criteria that appear frequently in the literature on tournament design: 1) Effectivity and efficacy; 2) Fairness; 3) Attractiveness; and 4) Strategy-proofness. Effectivity and efficacy are defined as the ability of a tournament to reveal the “true” (but hidden) ranking of participants. In contrast, fairness has no uniform definition (as will be discussed in Section 2.3.2); in our paper, it mainly refers to the principle of equal treatment. Attractiveness also has multiple interpretations, but they are always connected to the excitement that the tournament generates for its spectators. Finally, all participants should have an incentive to exert full effort and play according to their true strength under any possible scenario in the tournament. That is, the tournament design needs to be strategy-proof.
@@
Tournament design is a research field at the interface between operational research and (sports) economics. Operational research mainly contributes to tournament design by optimizing design choices over certain criteria and aiding decision-making. In sports economics, the main issue is typically how the tournament design may affect the behavior and efforts of the players, through theoretical or empirical research (Medcalfe,, 2024). Consequently, sports offer an excellent setting with skilled participants, high stakes, and transparent outcomes, to test various economic models and hypotheses (Palacios-Huerta,, 2023). There is a clear synergy between these two angles: sports economics research can bring new criteria and conditions to the table to be taken into account in optimization (or debunk myths that traditionally played a role in tournament design), and validate the decisions proposed by operational research.
@@ 2.1.1 Round robin
One of the basic competition formats is the (all-play-all) round robin tournament (Harary and Moser,, 1966). In a $k$-round robin ($k$RR) tournament, each participant faces each other $k$ times. Most domestic football leagues are $k$RR tournaments (Lasek and Gagolewski,, 2018); for example, the competitions in the top five leagues (England, France, Germany, Italy, Spain) are organized as a 2RR. Since 1RR is typically called “single round robin” and 2RR is called “double round robin”, we will use the notations SRR and DRR, respectively, for these tournament formats.
If not all rounds of a round robin tournament are played, by design (especially if there are many contestants), or because the tournament is aborted (see Section LABEL:Sec721), we call it an incomplete round robin tournament as teams do not meet all opponents an equal number of times. While this format dates back to at least Cochran, (1971), who discussed it in the context of bridge, it got more attention through the qualifying phase of the 2019-20 CONCACAF Nations League, which was contested by 34 teams in a single group, playing just 4 matches each (two home and two away). UEFA will also adopt an incomplete round robin structure for its club competitions from the 2024-25 season: the group stage will replaced by an incomplete round robin phase with 36 clubs that play against six or eight different opponents (UEFA,, 2023). Incomplete round robin tournaments can also emerge from unevenly combining all-play-all round robin tournaments between a subset of the opponents (e.g. teams in the same division) and bipartite round robin tournaments, where each team plays against each team from another group (e.g. teams in different conferences, or interleague play) as is common in North American sports.
@@ 2.1.2 Knockout
Another traditional competition format is the knockout tournament. In contrast to round robin tournaments, here a player generally faces only a subset of the other players. In a single knockout (SKO) tournament, players have only one chance of advancing to the next round: they are immediately eliminated if they lose.
In a double knockout (DKO) tournament, one loss is allowed in the tournament. This format is organized as two separate SKOs: the first SKO is called the main tournament, and the second SKO is called the consolation tournament. A player losing in the main tournament is moved to the consolation tournament. However, another loss in the consolation tournament eliminates the player. In essence, a player needs to lose twice in order to be knocked out. If the winner of the main round loses the final match against the winner of the consolation tournament, a rematch is played between them. A variant of this structure is used, for example, in judo, where the loser of the final in the main tournament fights against the winner of the consolation tournament for the second place. Thus, the winner will always have no losses, the silver medalist will have one loss, and the bronze medalist will have two losses.
In some knockout tournaments, the two players facing each other play more than one match. For instance, in the NBA playoffs, the opponents play as a “best-of-4”: the team that first reaches four wins advances to the next round, while the other team is eliminated. If a knockout tournament is repeated more than once among the same players, the SKO is said to be repeated (David,, 1959) or serial (Spieksma et al.,, 2022). The results of all subtournaments are aggregated to determine the final winner.
@@ 2.1.4 Multi-stage and hybrid tournaments
Tournaments can be played in multiple stages. Typically, only a subset of the players is allowed to proceed to the next stage, based on their performance in the previous stage. The FIFA World Cup, with a SRR group stage followed by a knock-out stage for the best teams, is a well-known example. In some tournaments, players accumulate points over different stages without being eliminated if they perform poorly in one stage. For instance, the Belgian domestic football championship has a regular season organized as a DRR. Then, the points accumulated by each team are halved, and the six best teams form a new group, the teams ranked 7–12 form a second group, and the remaining teams form the last group. Each group is again organized as a DRR; the winner of the top group becomes the league champion. Similar two-stage league formats are used in other European football leagues (Lasek and Gagolewski,, 2018). If the tournament formats differ in at least two stages, the tournament is commonly called hybrid.
@@ 2.2 Ranking
The ultimate aim of organizing a tournament is to rank the contestants by applying a ranking method. Section 2.2.1 introduces ranking in an individual tournament, Section 2.2.2 deals with the aggregation of rankings, and Section 2.2.3 gives a concise overview of global rankings, which rank the players on the basis of several tournaments, potentially all officially recognized matches played ever in the history of a sport.
@@ 2.2.1 Tournament ranking and tie-breaking
While various sophisticated ranking methods exist, the most widely used approach is to sum these scores. Naturally, this method does not take the quality of a win into account, i.e. matches played against a strong and a weak opponent yield the same number of points, which can only be fair if teams face the same opponents, preferably under the same conditions. This can be a problem especially in incomplete round robin and Swiss system tournaments, as will be discussed in Sections LABEL:Sec721 and LABEL:Sec723, respectively.
@@
Summing the scores resembles counting the number of opponents a player has beaten, which is the only ranking rule satisfying three appealing axioms (Rubinstein,, 1980): anonymity (the ranking method does not depend on the identity of the players), positive responsiveness (a loss does not result in a better rank than a win), and autonomous relative ranking (the relative ranking of two players is not influenced by a match where neither of them is involved in). Henriet, (1985) extends this characterization by allowing for ties, which are worth half of a win. While this method is used in some team sports where ties are allowed such as handball, the system that awards wins by three points has no theoretical background according to our knowledge.
@@
Furthermore, the (3,1,0) scoring system does not satisfy a well-known property in social choice theory called inversion (Chebotarev and Shamis,, 1998): the ranking should be reversed if all results are reversed. Consider a SRR with four players $i$, $j$, $k$, $\ell$ where $i$ won against $j$ and $\ell$, $j$ won against $\ell$, while the three matches of $k$ resulted in a draw. Using the notation ($=$, $>$) for the pairwise majority relation based on the aggregated score, the ranking according to the (3,1,0) method is $i>j>k>\ell$, and the ranking after all results are reversed is $\ell>j>k>i$. Naturally, the (2,1,0) scoring system satisfies inversion; in our example, the original ranking is $i>j=k>\ell$, and the ranking after all results are reversed is $\ell>k=j>i$.
@@
Often, a complete ranking of the players is desired. Therefore, ranking methods contain various tie-breaking rules to rank players with the same number of scores. In round robin tournaments, the most popular criteria are the number of matches won, head-to-head results, and superior goal difference (Csató, 2021c, , Chapter 1.3). The effects of choosing a specific hierarchy of tie-breaking rules are often non-negligible; for example, they influence the probability of heteronomous relative ranking, i.e. the relative ranking of two teams depends on the outcome of a match in which neither team was involved (Berker,, 2014).
@@
Knockout tournaments also require tie-breaking rules in some sports such as basketball, football, and handball. Tied matches are usually followed by extra time, and the final tie-breaking is usually a penalty shootout since the 1970s (Palacios-Huerta,, 2014). UEFA introduced the away goals rule in the 1965–66 European Cup Winners’ Cup for two-legged clashes. According to this principle, if both teams score the same number of goals in the two games, the team that scores more goals away will advance.
@@ 2.2.2 Aggregation of rankings
Some tournaments consist of individual contests, which have their own rankings. Examples include cycling (e.g. Tour de France), motorsport racing (e.g. Formula One), or winter sports (e.g. Biathlon World Cup). Here, a score is typically associated with each position in each contest, and they are added to obtain the final tournament ranking. A similar aggregation of rankings can be favourable even if the official rule is based on the “objective” finishing times: Ausloos, (2024) argues for ranking teams in cycling races by the cumulative sums of the places of their riders instead of their finishing times. However, the aggregation can lead to different paradoxes; Kaiser, (2019) presents some of them from Formula One.
Since rank aggregation is an extensively studied problem in social choice theory, we do not go into details. A recent summary is provided by Kondratev et al., (2023). The authors also find a one-parameter family of aggregation rules with geometric scores, which is the only one that satisfies two weak independence axioms: once an extremely strong or weak contestant is removed, the aggregate ranking is not allowed to change.
@@ 2.2.3 Global rankings
Gomes de Pinho Zanco et al., (2024) present a comprehensive stochastic analysis of the Elo algorithm in round robin tournaments, and investigate the effect of parameters $K$ and $\mu$ on its performance to obtain design guidelines. FIFA, for instance, after having received serious criticism in the academic literature (Lasek et al.,, 2016; Cea et al.,, 2020; Csató, 2021c, ; Kaminski,, 2022), adopted the Elo method of calculation in 2018 for their world ranking (FIFA,, 2018).
The popularity of Elo ratings is shown by the fact that the FIFA World Ranking has an alternative, the World Football Elo Ratings (https://www.eloratings.net/). The performance of football clubs is measured by at least two independent websites, http://clubelo.com/ and http://elofootball.com/. Nonetheless, some global rankings do not take the strength of opponents into account. For example, UEFA club coefficients are based on awarding points for wins and reaching certain stages in previous UEFA club competitions.
@@ 2.3 Design criteria
As mentioned in the Introduction, three main criteria have been identified to evaluate a tournament design: effectivity and efficacy, fairness, and attractiveness. They will be shortly discussed in Sections 2.3.12.3.3.
2.3.1 Effectivity and efficacy
The ability to select the best participant as the winner is one of the primary reasons why tournaments exist (Placek,, 2023). Glenn, (1960) calls this notion effectiveness, but it is also called predictive power (Ryvkin and Ortmann,, 2008; Vu and Shoham,, 2011). The ability of a tournament to rank all (and not just the best) contestants according to their true strength is called efficacy (Lasek and Gagolewski,, 2018; Sziklai et al.,, 2022). By far the most popular way of evaluating these properties is to use simulation: all matches of the tournament are simulated and the observed ranking is compared with the assumed true ranking of players. By aggregating the results over many simulation runs (i.e. Monte Carlo simulation), meaningful insights can be derived about how efficacious the tournament is. Efficacy is mostly studied from the perspective of the tournament format and the seeding rule.

### v2

@@ Abstract
Every sport needs rules. Tournament design refers to the rules that determine how a tournament, a series of games between a number of competitors, is organized. This study aims to provide an overview of the tournament design literature from the perspective of operational research. Three important design criteria are discussed: efficacy, fairness, and attractiveness. Our survey classifies the papers discussing these properties according to the main components of tournament design: format, seeding, draw, scheduling, and ranking. We also outline several open questions and promising directions for future research.
@@ 1 Introduction
Practicing sports is an important part of the life of many people all over the world as many health, psychological, and social benefits are associated with physical activity. Watching and following sports is even more popular than playing it. About 3 billion people watched the 2020 Summer Olympic Games (Statista, , 2023), while no less than 1.5 billion people tuned in for the 2022 FIFA World Cup final (FIFA, 2023b, ). Even tournament design aspects can be popular: the draw for the 2021/22 UEFA Champions League Round of 16 has attracted more than 1 million views on the YouTube channel of UEFA (Boczoń and Wilson, , 2023). Hence, “designing an optimal contest is both a matter of significant financial concern for the organizers, participating individuals, and teams, and a matter of consuming personal interest for millions of fans(Szymanski, , 2003, p. 1137).
@@
Every tournament is governed by a set of rules that determine the set of contestants, the format, the schedule, the ranking, and the prize allocation. Deciding on these rules is what we call tournament design. In the following, the contestants will be called players or teams interchangeably. The set of matches to be played is given by the tournament format. Seeding is typically used to guarantee that a strong contestant is not eliminated at the early stages of the tournament merely because it plays against an even stronger opponent. Subject to these and other constraints, the contestants can be allocated randomly, through a so-called draw. At first sight, the schedule, the order and timing of the matches, seems a purely operational issue (involving travel distance, referee and venue availability, etc.). However, it may have powerful effects on the fairness and the outcome of the tournament. Ranking aims to order the teams according to their performance in the tournament. Although performance can be directly measured in some sports (e.g. by finishing times), most sports allow only to rank the contestants by paired comparisons, games or matches. Prize allocation determines how prizes (e.g. promotion, relegation, qualification, broadcasting revenue) will be awarded to the contestants.
@@
Given the variety of tournament designs, the question arises on what basis they can be compared. We identify four different criteria that appear frequently in the literature on tournament design: 1) Efficacy; 2) Fairness; 3) Attractiveness; and 4) Strategy-proofness. Efficacy is defined as the ability of a tournament to reveal the “true” (but hidden) ranking of participants. In contrast, fairness has no uniform definition (see Section 2.3.2 and Baumer, (2024)); in our paper, it mainly refers to the principle of equal treatment. Attractiveness also has multiple interpretations, but they are always connected to the excitement that the tournament generates. Finally, all participants should have an incentive to exert full effort and play according to their true strength under any possible scenario in the tournament. That is, the tournament design needs to be strategy-proof.
Our focus on these criteria may suggest that sports federations can freely choose an optimal tournament design from their own perspective, i.e., a design showing an appropriate balance of theoretical properties. In practice, however, traditions and the nature of the sport may play an even larger role in this decision process. For instance, European football leagues are usually played in a round-robin format (Section 2.1.1), while tennis competitions are organized as a seeded knockout tournament (Section 2.1.2). Similarly, the Swiss system (Section 2.1.3) is mainly used in chess. Sports differ in how much time a player needs to recover between two matches: for instance, baseball teams play almost daily, while World Rugby recommends that at least four clear recovery days are provided after each match. The number of rest days required may depend on the frequency of changes allowed by the rules, too; ice hockey teams can play every second day mainly because they often change lines. Some sports plan their season in order avoid overlap with other popular sports (Pagels, , 2018), while travel costs may force tournaments to be organized in a short time span on one location (Nurmi et al., , 2014). Hence, any “good” tournament design depends on the particular sport, and operational research should take these constraints into account to improve decision making. However, in this paper – apart from a number of specific examples – we will maintain a more generic view.
@@
Tournament design is a research field at the interface between operational research and (sports) economics. Operational research mainly contributes to tournament design by optimizing design choices over certain criteria and aiding decision-making. In sports economics, the main issue is typically how the tournament design may affect the behavior and efforts of the players, through theoretical or empirical research (Medcalfe, , 2024; Palacios-Huerta, , 2023). There is a clear synergy between these two angles: sports economics research can bring new criteria and conditions to be taken into account in optimization (or debunk myths that traditionally played a role in tournament design), and validate the decisions proposed by operational research.
@@ 2.1.1 Round robin
One of the basic competition formats is the round robin tournament (Harary and Moser, , 1966). In a $k$-round robin ($k$RR) tournament, each participant faces each other $k$ times. Most domestic football leagues are $k$RR tournaments (Lasek and Gagolewski, , 2018); for example, the competitions in the top five leagues (England, France, Germany, Italy, Spain) are organized as a 2RR. Since 1RR is typically called “single round robin” and 2RR is called “double round robin”, we will use the notations SRR and DRR, respectively, for these formats.
If not all rounds of a round robin tournament are played, by design (especially if there are many contestants), or because the tournament is aborted (see Section LABEL:Sec721), we call it an incomplete round robin tournament. While this format dates back to at least Cochran, (1971), who discussed it in the context of bridge, it got more attention through the qualifying phase of the 2019/20 CONCACAF Nations League, which was contested by 34 teams in a single group, playing just 4 matches each (two home and two away). UEFA adopted an incomplete round robin structure for its club competitions from the 2024/25 season: the group stage is replaced by an incomplete round robin phase with 36 clubs that play against six or eight different opponents (UEFA, , 2023). Incomplete round robin tournaments can also emerge from unevenly combining round robin tournaments between a subset of the opponents (e.g. teams in the same division) and bipartite round robin tournaments, where each team plays against each team from another group (e.g. teams in different conferences, or interleague play) as is common in North American sports.
@@ 2.1.2 Knockout
Another traditional competition format is the knockout tournament. In contrast to round robin tournaments, here a player generally faces only a subset of the other players. In a single knockout (SKO) tournament, players are immediately eliminated if they lose.
In a double knockout (DKO) tournament, one loss is allowed in the tournament. This format is organized as two separate SKOs: the first SKO is called the main tournament, and the second SKO is called the consolation tournament. A player losing in the main tournament is moved to the consolation tournament. However, another loss in the consolation tournament eliminates the player. In essence, a player needs to lose twice in order to be knocked out. If the winner of the main round loses the final match against the winner of the consolation tournament, a rematch is played between them. The final phase of the Russian football cup is a double knockout tournament since the 2022/23 season. A variant of this structure is used, for example, in judo, where the loser of the final in the main tournament fights against the winner of the consolation tournament for the second place. Thus, the winner will always have no losses, the silver medalist will have one loss, and the bronze medalist will have two losses.
In some knockout tournaments, the two players facing each other play more than one match. For instance, in the NBA playoffs, the opponents play as a “best-of-7”: the team that first reaches four wins advances to the next round. If multiple knockout tournaments are played by the same players, the SKO is said to be repeated (David, , 1959) or serial (Lambers et al., , 2024). The results of all subtournaments are aggregated to determine the final winner.
@@ 2.1.4 Multi-stage and hybrid tournaments
Tournaments can be played in multiple stages. Typically, only a subset of the players is allowed to proceed to the next stage, based on their performance in the previous stage. If the tournament formats differ in at least two stages, the tournament is commonly called hybrid. The FIFA World Cup, with a SRR group stage followed by a knockout stage for the best teams, is a well-known example.
@@ 2.2 Ranking
The ultimate aim of organizing a tournament is to rank the contestants. Section 2.2.1 introduces ranking in an individual tournament, Section 2.2.2 deals with the aggregation of rankings, and Section 2.2.3 gives a concise overview of global rankings, which rank the players on the basis of several tournaments.
@@ 2.2.1 Tournament ranking and tie-breaking
While various sophisticated ranking methods exist, the most widely used approach is to sum these scores. Naturally, this method does not take the quality of a win into account, i.e. matches played against a strong and a weak opponent yield the same number of points. This can be a problem especially in incomplete round robin and Swiss system tournaments, as will be discussed in Sections LABEL:Sec721 and LABEL:Sec723, respectively.
@@
Summing the scores resembles counting the number of opponents a player has beaten, which is the only ranking rule satisfying three appealing axioms (Rubinstein, , 1980): anonymity (the ranking method does not depend on the identity of the players), positive responsiveness (a loss does not result in a better rank than a win), and autonomous relative ranking (the relative ranking of two players is not influenced by a match where neither of them is involved in). Henriet, (1985) extends this characterization by allowing for ties, which are worth half of a win. While this method is used in some team sports where ties are allowed such as handball, the (3,1,0) scoring system has no theoretical background according to our knowledge.
@@
Another well-known property from social choice theory is inversion (Chebotarev and Shamis, , 1998): the ranking is reversed if all results are reversed. Consider a SRR with four players $i$, $j$, $k$, $\ell$ where $i$ won against $j$ and $\ell$, $j$ won against $\ell$, while the three matches of $k$ resulted in a draw. This example shows that the (3,1,0) scoring system does not satisfy inversion. Indeed, using the notation ($=$, $>$) for the pairwise majority relation based on the aggregated score, the ranking according to the (3,1,0) method is $i>j>k>\ell$, and the ranking after all results are reversed is $\ell>j>k>i$. Naturally, the (2,1,0) scoring system satisfies inversion; in our example, the original ranking is $i>j=k>\ell$, and the ranking after all results are reversed is $\ell>k=j>i$.
@@
Often a complete ranking of the players is desired. Therefore, ranking methods contain various tie-breaking rules to rank players with the same score. In round robin tournaments, the most popular criteria are the number of matches won, head-to-head results, and superior goal difference (Csató, 2021c, , Chapter 1.3). The effects of choosing a specific hierarchy of tie-breaking rules are often non-negligible; for example, they influence the probability of heteronomous relative ranking, i.e. the relative ranking of two teams depends on the outcome of a match in which neither team was involved (Berker, , 2014).
@@
Knockout tournaments also require tie-breaking rules in some sports such as football and handball. Tied matches are usually followed by extra time, and the final tie-breaking is usually a penalty shootout since the 1970s (Palacios-Huerta, , 2014). UEFA introduced the away goals rule in the 1965/66 European Cup Winners’ Cup for two-legged clashes. According to this principle, if both teams score the same number of goals in the two games, the team that scores more goals away will advance.
@@ 2.2.2 Aggregation of rankings
Some tournaments consist of individual contests, which have their own rankings. Examples include cycling (e.g. Tour de France), motorsport racing (e.g. Formula One), or winter sports (e.g. Biathlon World Cup). Here, a score is typically associated with each position in each contest, and they are added to obtain the final tournament ranking. A similar aggregation of rankings can be favourable even if the official rule is based on the “objective” finishing times: Ausloos, (2024) argues for ranking teams in cycling races by the cumulative sums of the places of their riders instead of their finishing times. However, the aggregation can lead to different paradoxes; Kaiser, (2019) presents some of them from Formula One. Since rank aggregation is an extensively studied problem in social choice theory, we do not go into details; a recent summary is provided by Kondratev et al., (2023).
@@ 2.2.3 Global rankings
Naturally, the Elo method can be parameterized in different ways depending on the value of $K$, the definition of match outcome $S$, or the adjustment of $r_{i}$ due to the potential home advantage. Gomes de Pinho Zanco et al., (2024) present a comprehensive stochastic analysis of the Elo algorithm in round robin tournaments, and investigate the effect of parameters $K$ and $\mu$ on its performance to obtain design guidelines.
FIFA, for instance, after having received serious criticism in the academic literature (Lasek et al., , 2016; Cea et al., , 2020; Csató, 2021c, ; Kaminski, , 2022), adopted the Elo method for their world ranking (FIFA, , 2018). The popularity of Elo ratings is shown by the fact that the FIFA world ranking has an alternative, the World Football Elo Ratings (https://www.eloratings.net/), that takes home advantage and goal difference into account as suggested by Szczecinski and Roatis, (2022). Analogously, the performance of football clubs is measured by at least two independent websites, http://clubelo.com/ and http://elofootball.com/. Nonetheless, some global rankings do not take the strength of opponents into account. For example, UEFA club coefficients are based on awarding points for wins and reaching certain stages in previous UEFA club competitions (Dagaev and Rudyak, , 2019).
@@ 2.3 Design criteria
Sections 2.3.12.3.3 shortly discuss the three main criteria that have been identified to evaluate a tournament design: efficacy, fairness, and attractiveness.
2.3.1 Efficacy
The ability of a tournament to rank contestants according to their true strength is called efficacy (Lasek and Gagolewski, , 2018; Sziklai et al., , 2022). Efficacy with respect to only the best player is also known as effectivity (Glenn, , 1960) or predictive power (Ryvkin and Ortmann, , 2008; Vu and Shoham, , 2011).
By far the most popular way of evaluating these properties is to use simulation: all matches of the tournament are simulated and the observed ranking is compared with the assumed true ranking of players.By aggregating the results over many simulation runs (i.e. Monte Carlo simulation), meaningful insights can be derived about how efficacious the tournament is. Efficacy is mostly studied from the perspective of the tournament format and the seeding rule.